OpenStudy (anonymous):

f(x,y) = x/y, P=(2,1) and v= -1i -1j. Find the maximum rate of change of f at P. Find the (unit) direction vector in which the maximum rate of change occurs at P.

OpenStudy (anonymous):

Find the gradient of f(x,y), evaluate it at P, and then find the dot product with v

OpenStudy (anonymous):

Grad(f)=<1/y, -x/y^2> at P: <1, -2>

OpenStudy (anonymous):

so the dot product is: <1, -2>*<-1, -1> = -1 + 2 = 1

OpenStudy (anonymous):

oops. I skipped a step

OpenStudy (anonymous):

<\[<-1/\sqrt{2}, -1/\sqrt{2}>\]

OpenStudy (anonymous):

\[<1, -2>*<-1/\sqrt{2}, -1/\sqrt{2}>\], where * is a dot product

OpenStudy (anonymous):

so the final is \[1/\sqrt{2}\]

OpenStudy (anonymous):

I'm not asking for a directional derivative of f at P in the direction of V. I'm asking maximum rate of changing f at P and unit vector in which the maximum rate of changing occurs at P.

OpenStudy (anonymous):

ah, sorry. The magnitude of the gradient is the greatest rate of change

OpenStudy (anonymous):

\[<1, -2>=\]=\[\sqrt{5}\]

OpenStudy (anonymous):

where that is the magnitude (greatest rate)

OpenStudy (anonymous):

then the unit vector would be \[<1/\sqrt{5}, -2/\sqrt{5}>\]